The water in the water heater is heated at 80 ° C overnight and its average temperature must not fall below 55 ° C. How much water should be in the reservoir to store 25 kWh of thermal energy? The reservoir's own thermal capacity should be insignificant.
1
Expert's answer
2019-03-15T12:40:31-0400

1kWh = 3.6*106J

then 25kWh = 25*3.6*106 J = 9*107J

c(water) = 4200 J/(kg* 0C)

Q = cmΔT

-9*107J = 4200 J/(kg* 0C) *m* (550C - 800C)

m = 857 kg

Answer: 857 kg